OpenStudy (anonymous):

Operation on functions: f(x)=1/2x g(x)=4/x+12

terenzreignz (terenzreignz):

What operation?

OpenStudy (anonymous):

the suspense is killing me

OpenStudy (anonymous):

all,,xD

OpenStudy (anonymous):

sum diff product quo

terenzreignz (terenzreignz):

oh well, I have faith in you... maybe you've already done sum and difference? :D

OpenStudy (anonymous):

nope. hahahahahha.. Tanga ako bro. :)))))

OpenStudy (anonymous):

i never really liked math, xD

terenzreignz (terenzreignz):

Just to rid ourselves of ambiguities, \[f(x) = \frac12x\] \[g(x) = \frac4x + 12\] Are these correct?

OpenStudy (anonymous):

nope x +12 on the denomenator,

terenzreignz (terenzreignz):

Okay... \[\Large g(x) = \frac4{x+12}\]

OpenStudy (anonymous):

yup

terenzreignz (terenzreignz):

Okay, let's add them up... \[\Large \frac12x + \frac4{x+12}\]

OpenStudy (anonymous):

wait 1/2x x is also on the denominator

terenzreignz (terenzreignz):

You could have said that earlier :/ \[\Large \frac1{2x}+ \frac4{x+12}\] better now? :P

OpenStudy (anonymous):

yup, xD sorry didn't noticed that

terenzreignz (terenzreignz):

Okay, you know how to add fractions? Because it'd be much prettier if it had only one fraction bar, aight?

OpenStudy (anonymous):

yup,

terenzreignz (terenzreignz):

So can you add fractions?

OpenStudy (anonymous):

yup,

terenzreignz (terenzreignz):

Well then, add them :P

OpenStudy (anonymous):

but not with variables, xD

OpenStudy (anonymous):

can please just do the solution and explain it to me,, i still have to study chem ap and biotech, we have exams, T_T

terenzreignz (terenzreignz):

Fine, I'll do the addition, you do the subtraction (it's basically the same, anyway) Sound fair? :P

OpenStudy (anonymous):

agree

terenzreignz (terenzreignz):

The trick to adding or subtracting (dissimilar, or unlike denominators) fractions such as \[\Large {\frac1{2x}}+ {\frac4{x+12}}= \color{red}{\frac{\color{white}{xxxxxx}}{}}\] The denominator of the sum (difference) is simply the product of the denominators... \[\Large {\frac1{2x}}+ {\frac4{x+12}}= \color{red}{\frac{\color{white}{xxxxxx}}{2x(x+12)}}\] Catch me so far?

OpenStudy (anonymous):

yup.

terenzreignz (terenzreignz):

\[\Large {\frac1{\color{green}{2x}}}+ {\frac4{\color{blue}{x+12}}}= \color{red}{\frac{\color{white}{xxxxxx}}{\color{green}{2x}\color{blue}{(x+12)}}}\]

terenzreignz (terenzreignz):

In the numerator, you 'cross' it. What I mean is that the left numerator is multiplied to the right denominator... \[\Large {\frac1{\color{green}{2x}}}+ {\frac4{\color{blue}{x+12}}}= \color{red}{\frac{1\color{blue}{(x+12)}}{\color{green}{2x}\color{blue}{(x+12)}}}\] And the right numerator is multiplied to the left denominator \[\Large {\frac1{\color{green}{2x}}}+ {\frac4{\color{blue}{x+12}}}= \color{red}{\frac{1\color{blue}{(x+12)}+4(\color{green}{2x})}{\color{green}{2x}\color{blue}{(x+12)}}}\]

terenzreignz (terenzreignz):

Can you simplify this?

OpenStudy (anonymous):

\[\frac{ x+12 +8x}{2x ^{2}+24x }\]

terenzreignz (terenzreignz):

Well now, that's really good :P Except you forgot to combine these. \[\huge \frac{ \color{red}x+12 +\color{red}{8x}}{2x ^{2}+24x }\] YOU FAIL!!!!! haha just kidding Though seriously, combine those, and that's your answer :P

OpenStudy (anonymous):

\[\frac{ 9x+12 }{ 2x ^{2} }\]

OpenStudy (anonymous):

+24x

terenzreignz (terenzreignz):

Much better... Okay, general rule... regardless... \[\Large \frac{a}b \pm \frac{c}d = \frac{ad \pm bc}{bd}\]

OpenStudy (anonymous):

okay..i'm gonna study other subjects now, thanks by the way,,

terenzreignz (terenzreignz):

You haven't done product or quotient yet, though. Stay awhile, this'll take less than five minutes...

OpenStudy (anonymous):

i really need to go, i know how to answer the product and quo just needed the sum and difference.xD

OpenStudy (anonymous):

Please hold your peace, I meant no offense. I also didn't expect that much hostility from you. I'm in a good mood, I usually am, but if you really think the one you are having a comment war with on youtube is an idiot, it'd be best if you just let them be. This is the internet, if you expect every single one of the three hundred fifty million users to be friendly with you, you're gonna have a bad time.

OpenStudy (anonymous):

It looks like somebody has been working on the 'making comedy videos' thing... Nice to see that you appear to be doing fine :)

terenzreignz (terenzreignz):

someone's stalking me

OpenStudy (anonymous):

.xD hahaha

terenzreignz (terenzreignz):

ehh... Until we meet again... Derek... <insert evil laugh> Name's Teejay, for future reference I'll be sure not to underestimate you again...

OpenStudy (anonymous):

okaay, xD bye, thanks for the time though,